# Principles of Visual Tokens for Efficient Video Understanding

### v1

@@ Abstract
Video understanding has made huge strides in recent years, relying largely on the power of the transformer architecture. As this architecture is notoriously expensive and video is highly redundant, research into improving efficiency has become particularly relevant. This has led to many creative solutions, including token merging and token selection. While most methods succeed in reducing the cost of the model and maintaining accuracy, an interesting pattern arises: most methods do not outperform the random sampling baseline. In this paper we take a closer look at this phenomenon and make several observations. First, we develop an oracle for the value of tokens which exposes a clear Pareto distribution where most tokens have remarkably low value, and just a few carry most of the perceptual information. Second, we analyze why this oracle is extremely hard to learn, as it does not consistently coincide with visual cues. Third, we observe that easy videos need fewer tokens to maintain accuracy. We build on these and further insights to propose a lightweight video model we call LITE that can select a small number of tokens effectively, outperforming state-of-the-art and existing baselines across datasets (Kinetics400 and Something-Something-V2) in the challenging trade-off of computation (GFLOPs) vs accuracy.
Figure 1: Left: Given an input video, the proposed LITE method is able to select a subset (non-white patches) of discriminative tokens for any given computational budget. Right: LITE is lightweight and efficient, outperforming state-of-the-art in the trade-off between computational cost (GFLOPs) and accuracy. The plot shows the results on Something-Something-V2.
@@ 1 Introduction
Video understanding has made remarkable progress in the last few years, getting close to solving many standard benchmarks and tasks. This progress has largely hinged on the use of the transformer architecture [36], which is both extremely powerful as well as extremely computationally expensive. Transformers originated in the language field, where few tokens are needed to represent a concept, for example an action, as tokens roughly correspond to single words. This is not the case in the visual adaptation of transformers [1], where the number of tokens to represent an action is orders of magnitude larger, leading to an exponentially larger computational cost. This cost has a wide range of negative consequences: it limits the deployability of models, as they require expensive equipment to run, it burdens video understanding research, making turnaround slower, it is financially expensive and it has a big environmental footprint [16]. In particular, although the inference cost is small compared to training, over the lifetime of a model, the overall cost of inference is larger than the cost of training, as it is done many more times [28]. At the same time, video is notoriously redundant both in space (where background regions might be a large part of the scene) and time (with many frames being similar, even if they are downsampled from the original video. These two factors together pose a great opportunity opening a research theme on token reduction. This theme has led to many creative ideas on how to reduce the amount of tokens. One of the noticeable strands is token-merging, where similar tokens are grouped together such that only a representative of the group needs to be stored in memory [4, 9, 8, 14]. Other ideas have revolved around selecting tokens based on motion, such as the Motionformer [27]. These models can reduce the computational cost in GFLOPs considerably, while keeping accuracy on-par with the full model.
Crucially, we observe that most of these models make a trade-off between accuracy and computation that is similar or inferior to the simplest baseline of sampling tokens randomly (see Fig. 2). In this paper we take a closer look at this puzzling effect, and discover some underlying principles of the nature of visual tokens that are useful for designing faster, more efficient video models.
@@
Searching for the right tokens is like looking for a needle in a haystack. What does the oracle’s performance look like? Does it attend to cues that humans use, such as foreground objects? Furthermore, can we train a simple model to reproduce the oracle? We conducted a series of experiments to understand the nature of the tokens selected by the oracle. We assess how they compared to those selected by other methods, including: selecting the foreground, selecting the regions with more attention, etc. Surprisingly, all these strategies perform much worse than selecting tokens at random. We then propose a “selector” network, which we call LITE for LIghtweight Token Elector, which takes as input a token and approximates the value of the oracle. We observe that although the selector is extremely lightweight, it is enough to outperform previous work on token selection and merging, including the strong random-selection baseline. This selector is a simple multi-layer perceptron (MLP) which operates at the token level individually, and in our experiments no comparison methods give better performance.
@@
  • We incorporate these principles into a novel method, LITE that is architecture-agnostic and task-agnostic, and can be used as plug-and-play for other settings.
@@ 2 Related Work
Transformer Token Reduction. From all efficiency strategies, the most relevant to this work is reduction of tokens. Some methods drop tokens by predicting scores through end-to-end training [30, 44, 37]. For instance, DynamicViT [30] scores each token using a lightweight network and selects top tokens with Gumbel-softmax. A-ViT [44] computes halting scores from the token embedding’s first dimension to decide on pruning, while STTS [37] uses a differentiable Top-$K$ operator to predict scores and select top areas for reduced computation.
Additionally, some methods extract essential information from all tokens to reduce their number [26, 31, 20]. Strategies like $K$-centered search [26] are used for this purpose. TokenLearner [31] employs MLPs to learn a fixed number of new tokens, thus reducing the computational load from processing numerous tokens. LookupVit [20] uses compressed tokens for high computational processing, and other tokens for cheaper layers.
Another mainstream approach is merging redundant or similar tokens [4, 9, 8, 14]. The seminal ToMe [4] method merges similar tokens using bipartite matching algorithms. Building on ToMe, STA [9] developed a video-specific token dropping method where the first frame uses ToMe, and each subsequent frame is compared with the remaining tokens from the previous frame, retaining similar ones. Additionally, the vid-TLDR [8] work incorporates a saliency mask based on ToMe, enhancing the contribution of tokens in salient areas in subsequent learning. Moreover, the recent STTM [14] method extracts object cues, compresses token information within the same object, and add aggregated position information to compensate for the loss of motion cues. A special case of token reduction is frame selection, which has led to a series of interesting work including SMART [15], Sevila [45], Vila [39] and ATP [5].
@@ 3 Principles of Visual Tokens in Video
Datasets. We use two standard action recognition datasets: Kinetics-400 [7], sourced from YouTube, is a large-scale human action video dataset with 400 classes. Each class has a minimum of 400 action clips. Something-Something V2 (SS-V2) [18] focuses on fine-grained understanding of human hand gestures. Comprising 220,847 labeled video clips, it captures 174 individuals performing predefined actions with everyday items. This dataset poses a greater challenge for temporal modeling [33].
@@
  • Saliency: Finally we use a saliency detector CPFE [46] to compute the saliency at each frame. We then sample the tokens with highest saliency. Saliency turns out to not be better than random.
@@
Designing an oracle to predict the value of tokens. To identify which tokens play a decisive role in the classification task and to assign a score to each token based on their importance, we design an oracle to predict the value of a given token, using Grad-CAM  [32]. As mentioned before, we call this an oracle because it has access to privileged information – the true class label. However, this is indeed a tractable approximation of a “real oracle”, which would be to find the $k$-subset of tokens that minimizes the loss relative to the true label.
@@
Figure 6: Overview of LITE. At training time, video gradients are used to build the oracle, which the selector learns to replicate. At test time, the predicted value of the selector is used to keep the top most useful tokens, reducing the overall burden of the input to the classifier.
@@ 4.1 Review of Video Transformer
Video Transformer takes as input a clip of video $\mathcal{V}\in\mathbb{R}^{T\times H\times W\times 3}$, which contains $T$ RGB frames with the size of $H\times W$ sampled from the original video. Following the approach of ViT [10], Video Transformer flattens frames into $N$ 3D patches with size $\{p_{t},p_{h},p_{w}\}$, where $N=\lvert\frac{T}{p_{t}}\rvert\times\lvert\frac{H}{p_{h}}\rvert\times\lvert\frac{W}{p_{w}}\rvert$. Thus, Video Transformer converts a video to a sequence of non-overlapping patches.
In this paper, patch $n$ with temporal index $t$ is represented as $\mathbf{v}_{n}^{t}\in\mathbb{R}^{3p_{t}\times p_{h}\times p_{w}}$, where $n=1,...,N$ and $t=1,...,\lvert\frac{T}{p_{t}}\rvert$. The patch embedding vector $\mathbf{z}_{n}^{t}\in\mathbb{R}^{D}$ is obtained by means of a learnable linear projection and a separate addition of spatio-temporal positional embedding.
$\mathbf{z}_{n}^{t}=M\mathbf{v}_{n}^{t}+\mathbf{e}_{n}+\mathbf{e}^{t}$ (3)
$M$ is a learnable linear matrix and $e_{n}$ and $e^{t}$ are the separate spatial and temporal learnable positional embeddings. To model the appearance and motion in the video, each patch is processed with a stack of Transformer encoding blocks. In each block $\ell$, the patch embedding is updated by multi-head attention (Attention), LayerNorm (LN), and MLP processes:
$\hat{\mathbf{z}}_{n}^{t^{(\ell+1)}}=\operatorname{Attention}\left(\operatorname{LN}\left(\mathbf{z}_{n}^{t^{(\ell)}}\right)\right)+\mathbf{z}_{n}^{t^{(\ell)}}$ (4)
$\mathbf{z}_{n}^{t^{(\ell+1)}}=\operatorname{MLP}\left(\operatorname{LN}\left(\hat{\mathbf{z}}_{n}^{t^{(\ell+1)}}\right)\right)+\hat{\mathbf{z}}_{n}^{t^{(\ell+1)}}$ (5)
where $\ell=1,...,L-1$ and $L$ is the total number of blocks.
@@ 4.2 Learning to Select
Given the oracle concept, the goal is to design a lightweight selector that can predict the oracle’s value of each token. A necessary condition is that the model has to be lightweight, as it has to be significantly faster than using all tokens.
Assume a sequence of patch embeddings $\mathbf{p}\in\mathbb{R}^{N\times D}$ where $N$ is length of the patch embedding sequence and $D$ is the embedding dimension. To maintain simplicity and ensure the model remains lightweight, we used a three-layer fully connected (FC) network to predict token scores. Given that the scores in the oracle $s\in[0,1]$, we also incorporated a sigmoid function for better estimation:
$\mathbf{s}^{\prime}=\operatorname{Sigmoid}\left(\mathrm{FC}\left(\mathbf{p};\boldsymbol{\theta}\right)\right)\in\mathbb{R}^{N\times 1},$ (6)
where $\mathbf{p}\in\mathbb{R}^{N\times D}$ is the patch embedding, $\boldsymbol{\theta}$ is the network weights, and $\mathbf{s}^{\prime}\in\mathbb{R}^{N}$ is our estimated token score.
We train our selector in conjunction with the frozen patch embedding module from the backbone model, using binary cross-entropy loss between the selector’s and oracle’s token values. This is depicted in the yellow region of Fig. 6. During inference, we insert our selector between the patch embedding module and the first transformer block, utilizing all weights from the original backbone model except those of our selector. This is depicted in the green region of Fig. 6.
@@ 4.3 Adaptive Computational Budget
Principle 5 shows that not all videos or tasks need the same amount of computation, and easy ones need a smaller number of tokens. This is similar to the case of image recognition in human perception, where we perform certain tasks (such as detecting familiar objects) faster than harder tasks (such as classifying occluded objects) [29]. Here we explore the use of this principle as follows: we gauge the difficulty of each video in a very lightweight manner and use this information to compute an “adaptive budget”, which will adapt the computational budget and save computation on “easy” videos. This process is depicted as the orange box in Fig. 6 and it is referred to in experiments as LITE++, as it focuses on further reducing computation at a minimal cost.
@@ 4.4 Model Variants
Looking at the simplicity of LITE, one could think that learning to predict the oracle is particularly easy and straightforward. After all, if an MLP can learn this, potentially a more sophisticated model should outperform it. This section describes the series of variants evaluated, which might give additional insights into the nature of the token-selection problem as well as showcase that it is far from trivial. Results for all these variants are shown in Sec. 5.
Cleaning the oracle training data. The examples in Fig. 4 suggest that the oracle is somewhat noisy. Could there be some artifacts that we could correct to improve the performance of the oracle? Several cleaning strategies for the oracle were explored:
  • Edges. Since Grad-CAM often produces noisy high activation around image boundaries, we diminish the values on edges to reduce boundary artifacts.
  • Isolated peaks. The raw value of the oracle often has isolated peaks, which can misrepresent the true important areas. We remove small areas below a certain threshold.
  • Sharpening the distribution. The original distribution of the oracle is often smooth. We transform oracle values to amplify the differences between high and low values, making the distribution more concentrated around 0 or 1.
Including global information. The LITE model assesses the value of a token based solely on the information of one token. This is a simplification of the way tokens are used, and overlooks qualities such as diversity of information, relationship, etc. We experimented with adding a global branch to LITE , where instead of simply using the value of a token in isolation, we use several 3D convolutions to acquire the nearby context information. Then, a self-attention operation is employed to capture long-range dependencies and model interactions between tokens, allowing the network to integrate global information effectively.
Adding complexity to the selector. A 3-layer MLP is a fairly simple model, and potentially we should be able to do better with a more sophisticated model. Adding layers or using other architectures can potentially lead to better results, even at the cost of an increased amount of computation. We experimented with several aspects. We used more layers (up to 5) to add capacity to the MLP. We also tried other architectures and replace the MLP with single or multiple transformer blocks. Surprisingly, none of these variants impact the result of the selector. Results are shown in the Supplemental Material.
@@ 5 Experiments
The experimental validation on LITE starts with comparing current state-of-the-art models on the mainstream datasets Kinetics-400 and SS-V2. We then describe a series of studies to understand its behavior, including modifying architectural components, backbones, the oracle, etc. Experiment settings are the same as from the previous Sec. 3. Overall we observe that LITE outperforms previous state-of-the-art and that despite its simplicity is robust and effective.

### v2

@@ Abstract
Video understanding has made huge strides in recent years, relying largely on the power of transformers. As this architecture is notoriously expensive and video data is highly redundant, research into improving efficiency has become particularly relevant. Some creative solutions include token selection and merging. While most methods succeed in reducing the cost of the model and maintaining accuracy, an interesting pattern arises: most methods do not outperform the baseline of randomly discarding tokens. In this paper we take a closer look at this phenomenon and observe 5 principles of the nature of visual tokens. For example, we observe that the value of tokens follows a clear Pareto-distribution where most tokens have remarkably low value, and just a few carry most of the perceptual information. We build on these and further insights to propose a lightweight video model, LITE, that can select a small number of tokens effectively, outperforming state-of-the-art and existing baselines across datasets (Kinetics-400 and Something-Something-V2) in the challenging trade-off of computation (GFLOPs) vs accuracy. Experiments also show that LITE generalizes across datasets and even other tasks without the need for retraining.
Figure 1: Left: Given an input video (top row), the proposed LITE method is able to select a subset (non-white patches) of discriminative tokens for any given computational budget. Right: LITE is lightweight and efficient, achieving the optimal Pareto-front over state-of-the-art in the trade-off between computational cost (GFLOPs) and accuracy. The plot shows the results on Something-Something-V2.
@@ 1 Introduction
Video understanding has made remarkable progress in the last few years, getting close to solving many standard benchmarks and tasks [8, 21, 7]. This progress has largely hinged on the use of the transformer architecture [41], which is both extremely powerful as well as extremely computationally expensive. Transformers originated in the language field, where few tokens are needed to represent a concept, such as an action, as tokens roughly correspond to single words. This is not the case in the visual adaptation of transformers [1], where the number of tokens to represent an action is orders of magnitude larger, leading to an exponentially larger computational cost. This cost has a wide range of negative consequences: it limits the deployability of models, as they require expensive equipment to run, it burdens video understanding research, making turnaround slower, it is financially expensive and it has a big environmental footprint [18]. In particular, although the inference cost is small compared to training, over the lifetime of a model, the overall cost of inference is larger than the cost of training, as it is done many more times [32]. At the same time, video is notoriously redundant in space (where background regions might be a large part of the scene) and time (with many frames being similar, even if they are downsampled). These two factors together pose a great opportunity for token reduction and have led a research theme with many creative ideas. One of the noticeable strands is token merging, where similar tokens are grouped together such that only a representative of the group needs to be stored in memory [4, 11, 10, 16]. Another direction includes selecting tokens based on motion [31]. These models can reduce the computational cost in GFLOPs considerably, while keeping accuracy on-par with the full model. While this is promising, we observe that most token-reduction models make a trade-off between accuracy and computation that is similar or inferior to the simplest baseline of sampling tokens randomly (see Fig. 2). In this paper we take a closer look at this puzzling effect, to discover underlying principles of the nature of visual tokens. We leverage these principles to design faster, more efficient video models.
@@
Searching for the right tokens is like looking for a needle in a haystack. What does the oracle’s performance look like? Does it attend to cues that humans use, such as foreground objects? Furthermore, can we train a simple model to reproduce the information in the oracle? We conducted a series of experiments to understand the nature of the tokens selected by the oracle. We assess how they compared to those selected by other methods, including: selecting the foreground, selecting the regions with more attention, etc. Surprisingly, all these strategies perform worse than selecting tokens at random. We then propose a “selector” network, which we call LITE for LIghtweight Token Elector, which takes as input a token and approximates the value of the oracle. We observe that although the selector is extremely lightweight, it is enough to outperform previous work on token selection and merging, including the strong random-selection baseline. This selector is a simple multi-layer perceptron (MLP) which operates at the token level individually, and in our experiments no comparison methods give better performance.
@@
  • We incorporate these principles into a novel method, that can be used as plug-and-play for other settings.
@@
Note that this work focuses on visual tokens. This in practice means that our experiments use vision models and tasks (e.g., classification, zero-shot learning), instead of vision-language models (VLMs) and tasks (e.g., VQ&A, captioning). There are several reasons for this choice. First, while VLMs have shown excellent progress, much of their impressive nature is due to progress in language technology (e.g. LLMs and language embeddings) [48]. We find it urgent and essential to advance knowledge in the understanding of visual tokens, and how that will affect the development of future efficient and effective visual representations. Second, the distribution and properties of visual tokens from Vision Transformer (ViT) and vision-language tokens from VLMs are not necessarily similar. In particular, vision-language tokens in VLMs are often learned in a contrastive manner, to fit to the representation and distribution of language tokens which are different in nature [9]. Therefore we would not expect conclusions and properties of both families of tokens to be necessarily interconnected. We do experiment extensively with the zero-shot setting of datasets and tasks (which is one of the appealing properties of VLMs) and observe that the proposed model LITE successfully applies to novel datasets and tasks without the need for retraining.
@@ 2 Related Work
Transformer Token Reduction. From all efficiency strategies, the most relevant to this work is reduction of tokens. Some methods drop tokens by predicting scores through end-to-end training [34, 50, 42, 24]. For instance, DynamicViT [34] scores each token using a lightweight network and selects top tokens with Gumbel-softmax. A-ViT [50] computes halting scores from the token embedding’s first dimension to decide on pruning, while STTS [42] uses a differentiable Top-$K$ operator to predict scores and select top areas for reduced computation.
Additionally, some methods extract essential information from tokens to reduce their number [30, 35, 23]. Strategies like $K$-centered search [30] are used for this purpose. TokenLearner [35] employs MLPs to learn a fixed number of new tokens to reduce the computational load. LookupVit [23] uses compressed tokens for high computational processing, and other tokens for cheaper layers.
Another mainstream approach is merging redundant or similar tokens [4, 11, 10, 16]. The seminal ToMe [4] method merges similar tokens using bipartite matching algorithms. Building on ToMe, STA [11] developed a video-specific token dropping method where the first frame uses ToMe, and each subsequent frame is compared with the remaining tokens from the previous frame, retaining similar ones. Additionally, the vid-TLDR [10] work incorporates a saliency mask based on ToMe, enhancing the contribution of tokens in salient areas in subsequent learning. Moreover, the recent STTM [16] method extracts object cues, compresses token information within the same object, and add aggregated position information to compensate for the loss of motion cues. A special case of token reduction is frame selection, which has led to a series of interesting work including SMART [17], Sevila [51], Vila [44], and ATP [5].
@@ 3 Principles of Visual Tokens in Video
Datasets. We use three standard action recognition datasets and one spatial-temporal action detection dataset: Kinetics-400 (K400) [8], sourced from YouTube, is a large-scale human action video dataset with 400 classes. Each class has a minimum of 400 action clips. Something-Something V2 (SS-V2) [20] focuses on fine-grained understanding of human hand gestures. Comprising 220,847 labeled video clips, it captures 174 individuals performing predefined actions with everyday items. This dataset poses a greater challenge for temporal modeling [37]. UCF101 [38] is a dataset collected from YouTube, consisting of 13,320 video clips categorized into 101 different action classes. AVA Actions (AVA) [21] offers annotations for 80 atomic visual actions across 430 movie clips, each 15 minutes long. It totally provides 1.62M accurate spatial-temporal action labels.
@@
  • Saliency: We use a saliency detector CPFE [52] to compute the saliency at each frame and sample the tokens with highest saliency. It’s not better than random.
@@
Designing an oracle to predict the value of tokens. To identify which tokens play a decisive role in the classification task and to assign a score to each token based on their importance, we design an oracle to predict the value of a given token, using Grad-CAM  [36]. As mentioned before, we call this an oracle because it has access to privileged information—the true class label. However, this is indeed a tractable approximation of a “real oracle”, which would be to find the $k$-subset of tokens that minimizes the loss relative to the true label.
@@
Figure 6: Overview of LITE. At training time, video gradients are used to build the oracle, which the selector learns to replicate. At test time, the predicted value of the selector is used to keep the top most useful tokens, reducing the overall burden of the input to the classifier. Note that the dotted line connections are only performed during training.
@@ 4.1 Review of Video Transformer
The Video Transformer takes as input a video clip $\mathcal{V}\in\mathbb{R}^{T\times H\times W\times 3}$, and splits them into $N$ non-overlapping patches $\mathbf{v}_{n}^{t}\in\mathbb{R}^{p_{t}\times p_{h}\times p_{w}}$, where $n\in 1...n_{s}$, $n_{s}=\lvert\frac{H}{p_{h}}\rvert\times\lvert\frac{W}{p_{w}}\rvert$, $t\in 1...n_{t}$, $n_{t}=\lvert\frac{T}{p_{t}}\rvert$ and $N=n_{s}\times n_{t}$. These patches are converted into embeddings $\mathbf{p}_{n}^{t}\in\mathbb{R}^{D}$ by means of a learnable linear projection and a separate addition of spatial and temporal positional embedding, as in:
$\mathbf{p}_{n}^{t}=M\mathbf{v}_{n}^{t}+\mathbf{e}_{n}+\mathbf{e}^{t},$ (3)
where $M$ is a learnable linear matrix and $e_{n}$ and $e^{t}$ are the separate spatial and temporal learnable positional embeddings. These embeddings are then input to transformer blocks that perform self-attention. After that, these tokens are passed through a feedforward network, a 3-layer MLP that enables information exchange across channels.
@@ 4.2 Learning to Select
The goal at this stage is to consider the distribution of tokens $\mathbf{p}$ in the previous section and design a lightweight model that can predict the value of the oracle for each token. Based on this prediction, we can select the tokens with highest value to input to the transformer blocks. A necessary condition is that the model has to be lightweight, as it has to be significantly faster than using all tokens.
Assume a sequence of patch embeddings $\mathbf{p}$. To maintain simplicity and ensure the model remains lightweight, we used a three-layer fully connected multi-layer perceptron (MLP) network to predict token scores. Given that the scores $s$ from the oracle are in the range $[0,1]$, we incorporate a sigmoid function for better estimation:
$\mathbf{\hat{s}}=\operatorname{Sigmoid}\left(\mathrm{MLP}\left(\mathbf{p};\boldsymbol{\theta}\right)\right),$ (4)
where $\mathbf{p}$ is the patch embedding, $\boldsymbol{\theta}$ is the network weights, and $\mathbf{\hat{s}}$ is the sequence of all estimated tokens score.
We train our selector in conjunction with the frozen patch embedding module from the backbone model, using binary cross-entropy loss between the selector’s and oracle’s token values. It is shown in the yellow region of Fig. 6.
During inference, $\mathbf{\hat{s}}$ is computed, which contains the estimated score $\mathbf{\hat{s}}(n,t)$ for each embedding $\mathbf{p}_{n}^{t}$. For a given budget $b$ given as a percentage of the total number of tokens, we select the tokens with the highest scores.
This selector is inserted between the patch embedding module and the first transformer block, using all the weights of the original backbone model except those of our selector. This is depicted in the green region of Fig. 6.
@@ 4.3 Adaptive Computational Budget
Principle 5 shows that not all videos or tasks need the same amount of computation, and easy ones need a smaller number of tokens. This is similar to the case of image recognition in human perception, where we perform certain tasks faster than harder tasks [33]. Here we explore the use of this principle as follows: we gauge the difficulty of each video in a very lightweight manner and use this information to compute an “adaptive budget”, which will adapt the computational budget and save computation on “easy” videos. This process is depicted as the orange box in Fig. 6 and it is referred to in experiments as LITE++, as it focuses on further reducing computation at a minimal cost.
@@ 4.4 Model Variants
Given the simplicity of the MLP it would appear that performance could benefit from a more sophisticated model. We experiment extensively with three families of variants: 1) improving the oracle in different ways, 2) including global information that can allow the selector to choose tokens based on their joint value, and 3) adding complexity and capacity to the selector, for example using transformer blocks or adding layers. We report results for all these variants in the Supplementary Material and observe that none of them has an impact on the accuracy of the downstream task. This points to an interesting conclusion: the oracle is extremely hard to predict, and the MLP achieves a balance between accuracy and avoiding overfitting.
@@ 5 Experiments
The experimental validation on LITE first compares current state-of-the-art models on the classification task (Sec. 5.1), using the mainstream datasets Kinetics-400 and SS-V2 and observe that LITE achieves the pareto-front of the trade-off between computation and accuracy. We then evaluate the generalization capabilities of LITE on the zero-shot setting (Sec. 5.2) by testing a fixed selector (trained on K400) on other datasets (UCF101 [38] and SS-V2) without training or fine-tuning on them. We observe that accuracy is barely affected, showing that important tokens are consistent across domains. We take one step further and evaluate LITE on a new task (detection) and new dataset (AVA [21]) also without training. We again observe that mAP is barely affected while computation is reduced. We show extensive ablation experiments to understand the behavior of LITE (Sec. 5.4). Overall we observe that LITE outperforms previous state-of-the-art, generalizes well to new datasets and tasks and that despite its simplicity is robust and effective.